OpenStudy (anonymous):

Solve each of the quadratic equations below and describe what the solution(s) represent to the graph of each. Show your work to receive full credit. y = x^2 + 3x + 2 y = x^2 + 2x + 1

OpenStudy (tkhunny):

Factor, Factor...

OpenStudy (anonymous):

y=(x+3/2)^2 -1/4 or (y+1/4)= (x+3/2)^2 parabola again y = x^2 + 2x + 1 or y=(x+1)^2 again parabola for solution we have x^2 + 3x + 2 =x^2 + 2x + 1 or x=-1 and y =0 hence the two parabolas intersect at (-1,0)